OpenStudy (anonymous):

The slope of a function at any point (x,y) is e^x/e^x+1.The point (0,2ln2) is on the graph of f. write the equation of the tangent line to the graph of f at x=0

OpenStudy (anonymous):

evaluate the function at given point to get your y value. Then find derivative of function to get your slope. Then just plug in your points.

OpenStudy (anonymous):

i am unsure because the function given is slope

OpenStudy (anonymous):

not the original, and 0,2 is a random point on f not at the point of tangency

OpenStudy (anonymous):

woops, I see what you are saying let me re-read the question now. They gave you the slope of the function...not the function.

OpenStudy (anonymous):

i was thinking maybe if i take anti derivative i can get original function?

OpenStudy (anonymous):

and then maybe with original equation i can find y and then put into slope intercept?

OpenStudy (anonymous):

I'm actually a little confused on it now myself. I'll tell you what I was thinking and maybe it'll make sense to you... The point (0, 2ln2) is on the graph. so if we use point slope form y-2ln2=slope y-2ln2=(e^x)/(e^x +1) .... evaluate with x = 0 y-2ln2 = 1/2(x-0) y=(x/2) + 2ln2

OpenStudy (anonymous):

ya i got slope = 1/2 also

OpenStudy (anonymous):

but i think you cant use 2ln2 for equation of the line

OpenStudy (anonymous):

0,2ln2 isnt at the point of tangency

OpenStudy (anonymous):

although 0 is

OpenStudy (anonymous):

It's not at the point of tangency but it is a point on the original f(x). So I would assume you can get the original f(x) from it. When you evaluate f'(x) after getting the original equation you get (1/2).

OpenStudy (anonymous):

f'(x) at 0 you get (1/2)

OpenStudy (anonymous):

oh i c so it must be right then

OpenStudy (anonymous):

* wait 1/2 + 1

OpenStudy (anonymous):

y=(x/2) + 2ln2 is the original equation I got. Then when I took the derivative I got = 1/2+ 2(1/2) = 1/2 + 1 = 1/2 + 1 = 3/2

OpenStudy (anonymous):

From there I guess you would just plug that into point slope form to get the equation of the tangent line.

OpenStudy (anonymous):

wait why plug in again if we already got equation of original?

OpenStudy (anonymous):

shouldnt that be the answer?

OpenStudy (anonymous):

oh wait nvm

OpenStudy (anonymous):

got confused but i understnad lol

OpenStudy (anonymous):

i c maybe that will work i will try and let you know

OpenStudy (anonymous):

cool, let me know. It makes sense to me anyway, but I am not completely sure..... I'm going to break out my graphing calculator soon lol.

OpenStudy (anonymous):

off topic ,just realized maywather n cotto fight is today do u have a link?

OpenStudy (anonymous):

haha, nah I was actually just about to look for that myself as well.

OpenStudy (anonymous):

alright ill let you know if i find one